# Demonstrating CavePI: Autonomous Exploration of Underwater Caves by Semantic Guidance

### v1

@@ 8 Conclusion & Future Work
This work demonstrates CavePI, a portable and cost-effective AUV tailored for semantic-guided navigation in GPS-denied underwater environments. The platform’s key strength lies in extracting sparse semantic cues from noisy, low-light conditions through a fast and precise segmentation pipeline, enabling intelligent and adaptive navigation. The planning and control systems have been meticulously designed, rigorously tested, and fine-tuned within controlled laboratory settings to ensure robust and reliable performance. Comprehensive field deployments in diverse, challenging underwater environments, along with extensive digital twin simulations, have demonstrated CavePI’s reliability, adaptability, and operational efficiency. These results underscore its potential as a versatile and dependable tool for safe, long-term autonomous exploration of underwater caves. Future iterations will focus on optimizing system performance and expanding its capabilities to support a wider range of applications in marine robotics.

### v3

@@ 9 Conclusion & Future Work
Underwater cave exploration is important in hydrology for freshwater resource management, as well as in geological and archaeological studies. Since cave surveys are labor-intensive and hazardous, fully autonomous platforms can help significantly. This work demonstrates CavePI, a portable and cost-effective AUV tailored for semantic-guided navigation in GPS-denied underwater environments. The platform’s key strength lies in extracting sparse semantic cues from noisy, low-light conditions through a fast and precise segmentation pipeline, enabling intelligent and adaptive navigation. The planning and control systems have been meticulously designed, rigorously tested, and fine-tuned within controlled laboratory settings to ensure robust and reliable performance. Comprehensive field deployments in diverse, challenging underwater environments, along with extensive digital twin simulations, have demonstrated CavePI’s reliability, adaptability, and operational efficiency. These results underscore its potential as a versatile and dependable tool for safe, long-term autonomous exploration of underwater caves. Future iterations will optimize system performance and capabilities to support a wider range of applications for exploring any overhead structures (ship hulls, pipelines) without predefined markers. To achieve this, future work will focus on enabling zero-shot semantic understanding and active planning capabilities.
Acknowledgment
This research is supported in part by the U.S. National Science Foundation (NSF) grants #$2330416$, #$2024741$, and #$1943205$; and the University of Florida (UF) Research grant #$132763$. We are thankful to Dr. Nare Karapetyan, Ruo Chen, and David Blow for facilitating our field trials at Ginnie open-water springs and Blue Grotto. Furthermore, we acknowledge the help from Woodville Karst Plain Project (WKPP), El Centro Investigador del Sistema Acuífero de Quintana Roo A.C. (CINDAQ), Global Underwater Explorers (GUE), Ricardo Constantino, and Project Baseline over the years in providing access to challenging underwater caves and mentoring us in underwater cave mapping. The authors are also grateful for equipment support by Halcyon Dive Systems, Teledyne FLIR LLC, and KELDAN GmbH lights.